# Renormalization of the Quantum Stress Tensor Fluctuations and the Limits of Semiclassical Gravity

### v1

@@ Appendix A Four point function in 4d
Appendix B Renormalization details in a conformal 2d quantum field theory
The analysis of the previous section could be adapted directly to the computations performed here. However, because the two-point function of a massless scalar field $\phi$ is not well defined (due to infrared divergences) and due to special features associated with the conformal symmetry of the model, it is more convenient to carry out the renormalization of the energy–momentum tensor (and its fluctuations) directly in terms of the well-behaved and unambiguous correlation functions of $d\phi$ instead. Added interest of this section is that it starts with an explicit a self-contained, very concise, and explicit review of the covariant renormalization the expectation value $T_{\mu\nu}$ in 2d (more details of which can be found for instance in Fabbri:2005mw and references therein).
In 2d there is a Fock representation for each conformal coordinate system with the metric of the form (13) and two-point function in the associated vacuum is given by
@@ Appendix B Renormalization details in a conformal 2d quantum field theory
where $L$ is an infrared cutoff, and $\approx$ indicates that the expression is an approximation whose corrections play no role in the formulas that follow—those involving derivatives of the above correlation function—in the coincidence limit. The cut-off dependence disappears when considering the 2-point function of $d\phi$ instead, which is the one of interest in the definition of the energy momentum tensor for the conformal scalar field. Namely,
@@
On can regularize the expectation value of the energy momentum tensor by a point splitting procedure and the corresponding subtraction
@@
in a local neighbourhood of every point $p\in{\mathfs{M}}$. The classical expression of $T_{+-}$ vanishes identically due conformal invariance. The divergence $\nabla^{a}T_{ab}=g^{ca}(\partial_{c}T_{ab}-\Gamma^{d}_{ca}T_{db})$ in the conformal coordinates become $g^{+-}(\partial_{+}T_{-b}+\partial_{-}T_{+b})$ as the only non vanishing Christoffel symbols $\Gamma_{\pm\pm}^{\pm}$ drop due to $T_{+-}=0$. Using (31) we get
@@
Our task is to express $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}$ in terms of normal ordered quantities defined in an arbitrary representation so that we get an expression that is easy to compute with in some situations where a certain ‘vacuum state’ might capture a particular idealized physical situation. We then use (B.2) —written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$, namely
@@
where we used (B.2)—written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$ followed by equation (43). We can write the previous equation as follows
@@
$-\left[(-J(\xi_{12}^{X^{\pm}}))J(\xi_{34}^{X^{\pm}})+(\braket{1_{\pm}3_{\pm}}^{x}_{\psi}-J(\xi_{13}^{X^{\pm}}))J(\xi_{24}^{X^{\pm}})+(\braket{1_{\pm}4_{\pm}}^{x}_{\psi}-J(\xi_{14}^{X^{\pm}}))J(\xi_{23}^{X^{\pm}})\right$.{}
@@
$+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})]+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})]+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]+[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})][\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})]+[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})][\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]+[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})][\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]{}$
Writing the last line separately
@@
Finally, using (32) we get the following expression of the renormalized expectation value of the square of the energy momentum tensor Problema de signo creo en lo que sigue:
@@
- $\frac{\hbar}{2\pi}\braket{\psi|:T_{\pm\pm}:|\psi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)+\frac{\hbar^{2}}{48\pi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)^{2},{}$
whose covariance is granted by construction from the definition (44). The same method allows to write the $n$-point correlation functions for arbitrary $n$ in a close manner.
B.3 Calculation of $\braket{\psi|T_{++}T_{--}|\psi}_{\rm ren}$
@@
B.4 Operator product expansion
@@ Appendix C Relation to 4d
where we used ansatz $\Phi(u,v,\theta,\varphi)=\phi_{\ell}(u,v)Y_{\ell m}(\theta,\varphi)/r$ with $u=t-r_{*}$, $v=t+r_{*}$, $r_{*}=r+2M\log(r/(2M)-1)$ and the radial coordinate should be taken as an implicit function of the null coordinates, $r=r(u,v)$. At the horizon $r=2M$ the previous equation becomes the 2d conformally invariant field equation of a massless scalar field for all the modes, namely

### v2

@@ Appendix B Four point function in 4d
Appendix C Renormalization details in a conformal 2d quantum field theory
The analysis of the previous section could be adapted directly to the computations performed here. However, because the two-point function of a massless scalar field $\phi$ is not well defined (due to infrared divergences) and due to special features associated with the conformal symmetry of the model, it is more convenient to carry out the renormalization of the energy–momentum tensor (and its fluctuations) directly in terms of the well-behaved and unambiguous correlation functions of $d\phi$ instead. Added interest of this section is that it starts with a self-contained, concise, and explicit review of the covariant renormalization the expectation value $T_{\mu\nu}$ in 2d (more details of which can be found for instance in Fabbri:2005mw and references therein).
In 2d there is a Fock representation for each conformal coordinate system with the metric of the form (13) with the field operator written in the canonical form
$\phi(x)=\int_{0}^{\infty}\frac{d\omega}{4\pi\omega}\left(a^{+}_{\omega}e^{-i\omega x^{+}}+a^{-}_{\omega}e^{-i\omega x^{-}}+h.c.\right),$ (34)
with $a^{\pm}_{\omega}$ the annihilation operators for right and left moving modes with frequency $\omega$. The two-point function in the associated vacuum is IR divergent and thus ambiguous, yet it can be regularized by introducing an infrared cut-off scale $L$ and write
@@ Appendix C Renormalization details in a conformal 2d quantum field theory
where $\approx$ indicates that the expression is an approximation whose corrections play no role in the formulas that follow—those involving derivatives of the above correlation function—in the coincidence limit. The cut-off dependence disappears when considering the 2-point function of $d\phi$ instead, which is the one of interest in the definition of the energy momentum tensor for the conformal scalar field. Namely,
@@
For further use we will denote respectively $\Delta^{\pm}\equiv x^{\pm}_{1}-x^{\pm}_{2}$ and will often simply use just $\Delta$ when the context is clear. On can regularize the expectation value of the energy momentum tensor by a point splitting procedure and the corresponding subtraction
@@
in a local neighbourhood of every point $p\in{\mathfs{M}}$, and where $R$ is the scalar curvature (see (45) below). The classical expression of $T_{+-}$ vanishes identically due to conformal invariance. The divergence $\nabla^{a}T_{ab}=g^{ca}(\partial_{c}T_{ab}-\Gamma^{d}_{ca}T_{db})$ in the conformal coordinates become $g^{+-}(\partial_{+}T_{-b}+\partial_{-}T_{+b})$ as the only non vanishing Christoffel symbols $\Gamma_{\pm\pm}^{\pm}$ drop out due to $T_{+-}=0$. Using (40) we get
@@
Our task is to express $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}$ in terms of normal ordered quantities defined in an arbitrary representation so that we get an expression that is easy to compute with in some situations where a certain ‘vacuum state’ might capture a particular idealized physical situation. We then use (C.2) —written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$, namely
@@
where we used (C.2)—written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$ followed by equation (52). We can write the previous equation as follows
@@
$-\left[(\braket{1_{\pm}2_{\pm}}^{x}_{\psi}-J(\xi_{12}^{X^{\pm}}))J(\xi_{34}^{X^{\pm}})+(\braket{1_{\pm}3_{\pm}}^{x}_{\psi}-J(\xi_{13}^{X^{\pm}}))J(\xi_{24}^{X^{\pm}})+(\braket{1_{\pm}4_{\pm}}^{x}_{\psi}-J(\xi_{14}^{X^{\pm}}))J(\xi_{23}^{X^{\pm}})\right$.{}
@@
$+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})]+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})]+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]{}$
Writing the last equality separately
@@
Finally, using (41) we get the following expression of the renormalized expectation value of the square of the energy momentum tensor:
@@
$+$ $\frac{\hbar}{2\pi}\braket{\psi|:T_{\pm\pm}:|\psi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)+\frac{\hbar^{2}}{48\pi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)^{2},{}$
whose covariance is granted by construction from the definition (53). The same method allows to write the $n$-point correlation functions for arbitrary $n$ in a closed manner.
C.3 Calculation of $\braket{\psi|T_{++}T_{--}|\psi}_{\rm ren}$
@@
C.4 Operator product expansion
@@ Appendix D Relation to 4d
where we used the ansatz $\Phi(u,v,\theta,\varphi)=\phi_{\ell}(u,v)Y_{\ell m}(\theta,\varphi)/r$ with $u=t-r_{*}$, $v=t+r_{*}$, $r_{*}=r+2M\log(r/(2M)-1)$ and the radial coordinate should be taken as an implicit function of the null coordinates, $r=r(u,v)$. At the horizon $r=2M$ the previous equation becomes the 2d conformally invariant field equation of a massless scalar field for all the modes, namely
@@
Appendix E Squeezing is a relative notion
It is well known that in quantum filed theory and for generic space-times there is no canonical notion of a vacuum state. In certain situations (for instance when the space-time possesses a time translation symmetry reflected by the existence of a time-like killing field) one can use the associated notion of positive frequency to characterize the modes of the field, and use that notion to define a preferred vacuum state. Concretely, in the case of a scalar field (used here for illustration) one defines the field operator acting in the Fock representation via the sum
$\phi(x)=\sum_{i}a_{i}f_{i}(x)+a^{\dagger}_{i}\bar{f}_{i}.$ (73)
where $\{f_{i}\}$ is a basis of positive frequency solutions of the field equations and the coefficients $a_{i}$ and $a_{i}^{\dagger}$ are annihilation and creation operators operators. The associated vacuum state $|0\rangle$ is defined by the condition $a_{i}|0\rangle=0$. One can define a related state $|{\tilde{0}}\rangle$ that is considered as “squezzed” with respect to the former by considering a Bogoliubov transformation $b_{j}=\alpha_{j}^{i}a_{i}+\beta^{i}_{j}a^{\dagger}_{i}$ with $\bm{\alpha}^{\dagger}\bm{\alpha}-\bm{\beta}^{\dagger}\bm{\beta}={\bf 1}$ and the demand that $b_{i}|{\tilde{0}}\rangle=0$. It is clear that squezzing is then a relative notion, with the state $|{\tilde{0}}\rangle=0$ being squezzed w. r. t. $|0\rangle$ and the latter being squezzed w. r. t. the former. Both, the vaccum states and the related squezzed states have the property that the expectation value of the field vanishes. The renormalized energy momentum tensor in such states needs not vanish (as is well known say for the type of states that occur in the study of black hole evaporation) illustrating the fact that, generically, such states differ locally from the so called “freely falling vacuum” ( that is the state constructed using local Riemann coordinates as we have explicitly done here 2d Fabbri:2005mw ). It is for those cases that according to our explorations the quantum uncertainties of the energy momentum tensor are as large as the energy momentum tensor expectation value itself, and thus would fail, according to our proposed criterion, to provide suitable states for semi-classical gravity.